OpenStudy (anonymous):

I think of a number, multiply it by 4 subtract 6 and then divide by 5. If the answer is 10, what is the original number?

OpenStudy (anonymous):

19

OpenStudy (anonymous):

:)

jimthompson5910 (jim_thompson5910):

"I think of a number", we'll call it n for now "multiply it by 4" so we now have n x 4 or 4n "subtract 6" we now have 4n - 6 "then divide by 5" and we get \(\Large \frac{4n-6}{5}\)

jimthompson5910 (jim_thompson5910):

"If the answer is 10" means we set it equal to 10 and we can solve for n \[\Large \frac{4n-6}{5} = 10\] \[\Large 4n-6 = 10*5\] \[\Large 4n-6 = 50\] \[\Large 4n = 50+6\] \[\Large 4n = 56\] \[\Large n = \frac{56}{4}\] \[\Large n = 14\]

jimthompson5910 (jim_thompson5910):

So the original number is 14